# A Syntax-Injected Approach for Faster and More Accurate Sentiment Analysis

### v1

@@ 3.3 Results
Table 8: Comparison between SELSP, Stanza and VADER in accuracy (percent) in the polarity prediction task in the Spanish Rest-Mex training dataset (10 000 reviews) across dictionaries based on three polarity labels "neutral", "positive" and "negative"
Finally, we compare our models to a supervised Transformer-based model on the Rest-Mex Spanish dataset. Results are shown in Table  9. Note that, as explained in Section 3.2, this comparison is performed on a different split (the Rest-Mex test set) and we perform five-class classification, rather than three class, so results are not directly comparable to those in Table 8. The table shows a clear advantage of the Transformer-based model in terms of accuracy. This result is not surprising, since supervised models are known to excel when trained on a corpus with similar texts as the test data (Vilares et al., 2017), as is the case here. The use case for unsupervised lexicon-based models, like the rest of the systems considered in this paper, is when no in-distribution data is available to train a supervised model, a common situation in production.
@@
4 Conclusion and Discussion
In this study, we have addressed the speed bottleneck problem of sentiment analysis systems that use syntactic parsing. For this purpose, we trained a model for syntactic parsing as sequence labeling (SELSP) based on a pre-trained language model DistilBERT within the CoDeLIn formalism for sequence labeling. We then applied it to sentiment analysis on Spanish and English datasets, by using its output parse trees to compute sentence polarities from sentiment words via syntax-based rules.
@@ 4 Conclusion and Discussion
To put our results into context, we also ran a comparison with a supervised system based on variants of the Transformer-based model RoBERTa, which achieved state-of-the-art results at the Rest-Mex 2023 shared task. The supervised system achieved worse speed, but better accuracy, than our proposal in our comparison. This accuracy advantage is expected, as supervised systems trained or fine-tuned with in-distribution data (in this case, with more than 250K instances of reviews of the same type it is being evaluated with) tend to outperform unsupervised, rule-based systems as they can adapt specifically to the characteristics of the training data. However, this trend has been shown to often invert when one does not have in-domain training data, or even when the data is in-domain but from a different source or time period (see (Vilares et al., 2017)). For this reason, unsupervised approaches like the ones we focus on here are still popular in industry applications, where it is often not possible (or very costly) to obtain large amount of data that closely resembles production data.
@@
We noticed a big difference in accuracy across datasets (OpeNER vs. TripAdvisor reviews). This difference is not specific to our proposed system, but also common to all baselines, and it may be related with the length of the reviews (see § 3.1). We believe that the performance of SELSP in processing short reviews could be improved in the future by using specialized large language models that were pre-trained on short messages like tweets, such as BERTweet Nguyen et al. (2020). As our model depends on the syntactic relation between sentiment words and polarity shifters like negation, improving these rules can also lead to an improvement in accuracy.
@@ 5 Limitations
6 Declarations
6.1 Availability of data and materials
The authors confirm that the data supporting the findings of this study are available within the article and its supplementary materials.
@@ 6.1 Availability of data and materials
6.2 Funding
We acknowledge the European Research Council (ERC), which has funded this research under the Horizon Europe research and innovation programme (SALSA, grant agreement No 101100615), SCANNER-UDC (PID2020-113230RB-C21) funded by MICIU/AEI/10.13039/501100011033, Xunta de Galicia (ED431C 2020/11), and Centro de Investigación de Galicia “CITIC”, funded by Xunta de Galicia and the European Union (ERDF - Galicia 2014–2020 Program), by grant ED431G 2019/01.
6.3 Authors’ contributions
@@ Acknowledgements
We acknowledge the Rest-Mex Shared Task 2023 organizers Álvarez Carmona et al. (2023) for providing us the possibility to evaluate the results in Table  9, given that the test dataset is not accessible to the general public.
Appendix A Lists of English Negation Words and Modifiers
@@
Appendix B Conversion Rules
@@
from 0 to 2 = "Negative"
from 2 to 3 = "Neutral"
from 3 to 5 = "Positive"
@@
compound score <= -0.05 = "Negative"
(compound score > -0.05) and
(compound score < 0.05) = "Neutral"
compound score >= 0.05 = "Positive"

### v2

@@ Introduction
Table 8: Comparison between SELSP, Stanza and VADER in accuracy (percent) in the polarity prediction task in the Spanish Rest-Mex training dataset (10 000 reviews) across dictionaries based on three polarity labels ”neutral”, ”positive” and ”negative”
Finally, we compare our models to a supervised Transformer-based model on the Rest-Mex Spanish dataset. Results are shown in Table  9. Note that, as explained in Section “Methodology”, this comparison is performed on a different split (the Rest-Mex test set) and we perform five-class classification, rather than three class, so results are not directly comparable to those in Table 8. The table shows a clear advantage of the Transformer-based model in terms of accuracy. This result is not surprising, since supervised models are known to excel when trained on a corpus with similar texts as the test data (Vilares et al.,, 2017), as is the case here. The use case for unsupervised lexicon-based models, like the rest of the systems considered in this paper, is when no in-distribution data is available to train a supervised model, a common situation in production.
@@
Conclusion
In this study, we have addressed the speed bottleneck problem of sentiment analysis systems that use syntactic parsing. For this purpose, we trained a model for syntactic parsing as sequence labeling (SELSP) based on a pre-trained language model DistilBERT within the CoDeLIn formalism for sequence labeling. We then applied it to sentiment analysis on Spanish and English datasets, by using its output parse trees to compute sentence polarities from sentiment words through a sentiment analysis engine which relies on syntax-based rules.
@@ Conclusion
To put our results into context, we also ran a comparison with a supervised system based on variants of the Transformer-based model RoBERTa, which achieved state-of-the-art results at the Rest-Mex 2023 shared task. The supervised system achieved worse speed, but better accuracy, than our proposal in our comparison. This accuracy advantage is expected, as supervised systems trained or fine-tuned with in-distribution data (in this case, with more than 250K instances of reviews of the same type it is being evaluated with) tend to outperform unsupervised, rule-based systems as they can adapt specifically to the characteristics of the training data. However, this trend has been shown to often invert when one does not have in-domain training data, or even when the data is in-domain but from a different source or time period (see (Vilares et al.,, 2017)). For this reason, unsupervised approaches like the ones we focus on here are still popular in industry applications, where it is often not possible (or very costly) to obtain large amount of data that closely resembles production data.
@@
We noticed a big difference in accuracy across datasets (OpeNER vs. TripAdvisor reviews). This difference is not specific to our proposed system, but also common to all baselines, and it may be related with the length of the reviews (see § 3.1). We believe that the performance of SELSP in processing short reviews could be improved in the future by using specialized large language models that were pre-trained on short messages like tweets, such as BERTweet (Nguyen et al.,, 2020). As our model depends on the syntactic relation between sentiment words and polarity shifters like negation, improving these rules can also lead to an improvement in accuracy.
@@
Declarations
Availability of data and materials
The authors confirm that the datasets (OpeNERen, OpeNERes, UD_English-EWT and UD_Spanish-AnCora), code and sentiment dictionaries that support the findings of this study are available in the Zenodo repository  [^9] https://doi.org/10.5281/zenodo.15323755. However, restrictions apply to the availability of the Rest-Mex 2023 dataset, which was used under license for this study. With the permission of Rest-Mex 2023, dataset can be obtained from  [^10] https://sites.google.com/cimat.mx/rest-mex2023.
@@
Funding
We acknowledge the European Research Council (ERC), which has funded this research under the Horizon Europe research and innovation programme (SALSA, grant agreement No 101100615), SCANNER-UDC (PID2020-113230RB-C21) funded by MICIU/AEI/10.13039/501100011033, LATCHING (PID2023-147129OB-C21) funded by MICIU/AEI/10.13039/501100011033 and ERDF (EU), Ministry for Digital Transformation and Civil Service and “NextGenerationEU” PRTR under grant TSI-100925-2023-1, Xunta de Galicia (ED431C 2024/02), and Galician Research Center “CITIC”, funded by Xunta de Galicia through the collaboration agreement between the Consellería de Cultura, Educación, Formación Profesional e Universidades and the Galician universities for the reinforcement of the research centres of the Galician University System (CIGUS).
Authors’ contributions
@@ Acknowledgements
We acknowledge the Rest-Mex Shared Task 2023 organizers Álvarez-Carmona et al., (2023) for providing us the possibility to evaluate the results in Table 9, given that the test dataset is not accessible to the general public.
Appendix A
Lists of English Negation Words and Modifiers
@@
Appendix B
Conversion Rules
@@
from 0 to 2 = ”Negative”
from 2 to 3 = ”Neutral”
from 3 to 5 = ”Positive”
@@
[…]